OpenStudy (anonymous):

Write an equation of a line perpendicular to y = 1/4 x + 4 in slope-intercept form that passes through the point (-2, 6).

OpenStudy (anonymous):

a. y= 1/4 x - 4 b. y= 1/4 x + 6 c. y= -4x - 2 d. y= -4x +1/4

OpenStudy (anonymous):